# Chen, Simchi-Levi, and Wang

### v1

@@ 6 Estimation with Perfect Information about the Control Arm
Since we are implementing a fixed policy $\pi^{\mathfrak{t}}$ during the experimental periods, it is logical that the asymptotic variance $\Sigma_{\mathrm{PI}}$ and $\sigma^{2}_{\mathrm{PI}}(s^{1})$ exhibit similar structures with $\Sigma_{\mathrm{MB}}$ and $\sigma^{2}_{\mathrm{MB}}(s^{1})$ from the model-based estimator under a fixed policy discussed in Section 3. Unlike the approach in that section, which requires estimating transitions for every state, here we only need to estimate the transitions and rewards associated with the crucial state $s^{1}$. Recalling Corollary 3.5, the asymptotic variance $\sigma^{2}_{\textrm{MB}}(s^{1})$ is the sum of variances associated with all states. With PI, $\sigma^{2}_{\textrm{PI}}(s^{1})$ only keeps the variance linked to the crucial state $s^{1}$, getting rid of all the other terms in $\sigma^{2}_{\textrm{MB}}(s^{1})$.
This reduction aligns with our expectations, given the complete information on the transitions and rewards for non-crucial states now at our disposal.
@@ 8 Simulations
We revisit the example of customer lifetime value (Example 1.1), with a synthetic transition matrix detailed in Appendix. We assume that under the control policy, when the customer makes a purchase, the company will receive a net income of $30$ while the income in all other states is 00. We consider the following two types of treatment:
@@
    SST. The company sends a coupon when the customer is at highest recency (recency=5).
@@
    Local Treatment. The company starts posting advertisements whenever the customer’s recency is high (recency$>$2).
For each type of treatment, we consider $5$ candidate coupon values (advertising cost), i.e. $n=5$ test arms (the coupon value (advertising cost) is 00 for the control arm). We then run experiments for each treatment type, aiming to identify the optimal coupon value (advertising cost) and report their ATE. In each experiment, we run all three estimators over 1,000 independent trajectories and record their empirical mean and standard error. The experiment horizon is adjusted so that the IS estimator’s standard error is smaller than its empirical mean.
@@
Unbiased Estimation Across All Estimators: The empirical means for all three estimators in both experiments demonstrate negligible bias. This aligns with our theoretical prediction of asymptotic normality for all three estimators.
@@
IS and PI work: In both examples, we observe a significant reduction in variance—by at least half—through the application of IS and PI. This corroborates the effectiveness of IS and PI in overcoming the lower bound of A/B testing by utilizing the local structure. Moreover, PI consistently outperforms IS in both examples, confirming its capability in further variance reduction.
@@
Local Treatment is Harder than SST: While ATEs of each experiment are close, the local treatment required a substantially longer horizon for the standard error of IS to decay compared to SST. This discrepancy arises because the SST allows information sharing for more states compared to local treatment. This comparison clearly illustrates that the efficacy of IS increases with the number of shared states.
@@ 9.2 Bias-Variance Trade-off
With general treatment, the seminal work by Farias et al. (2022) proposes a Difference-in-Q (DQ) estimator that trades a small bias for a striking variance reduction in average reward setting. In a follow-up work (Farias et al. 2023), they extend the idea of DQ to other reward settings including discounted reward used in this paper. In particular, we adapt their DQ theory to obtain the following estimator

### v2

@@ 6 Estimation with Perfect Information about the Control Arm
Since we are implementing a fixed policy $\pi^{\mathfrak{t}}$ during the experimental periods, it is logical that the asymptotic variance $\Sigma_{\mathrm{PI}}$ and $\sigma^{2}_{\mathrm{PI}}(s^{1})$ exhibit similar structures with $\Sigma_{\mathrm{MB}}$ and $\sigma^{2}_{\mathrm{MB}}(s^{1})$ from the model-based estimator under a fixed policy discussed in Section 3. Unlike the approach in that section, which requires estimating transitions for every state, here we only need to estimate the transitions and rewards associated with the crucial state $s^{1}$. Recalling Corollary 3.5, the asymptotic variance $\sigma^{2}_{\textrm{MB}}(s^{1})$ is the sum of variances associated with all states. With PI, $\sigma^{2}_{\textrm{PI}}(s^{1})$ only keeps the variance linked to the crucial state $s^{1}$, getting rid of all the other terms in $\sigma^{2}_{\textrm{MB}}(s^{1})$. This reduction aligns with our expectations, given the complete information on the transitions and rewards for non-crucial states now at our disposal.
@@ 8 Simulations
8.1 Synthetic Data: Customer Lifetime Value
We revisit the example of customer lifetime value (Example 1.1), with a synthetic transition matrix detailed in Appendix 19.1. We assume that under the control policy, when the customer makes a purchase, the company will receive a net income of $30$ while the income in all other states is 00. Specifically, the transition matrix and the reward under the control policy are set as follows.
${\bm{P}}^{\mathfrak{c}}=\begin{bmatrix}0.6&0.4&&&\\ 0.5&&0.5&&\\ 0.4&&&0.6&\\ 0.3&&&&0.7\\ 0.1&&&&0.9\end{bmatrix}\,,\qquad{\bm{r}}^{\mathfrak{c}}=\begin{bmatrix}30\\ 0\\ 0\\ 0\\ 0\end{bmatrix}\,.$
We consider the following two types of treatment:
@@ 8.1 Synthetic Data: Customer Lifetime Value
    SST. The company considers to offer coupons when the customer reaches the highest recency (recency = 5). We evaluate five different candidate coupon values, each associated with a different treatment effect. These treatments lead to a decrease in reward at state $s^{5}$ by values of $2,2.5,3,3.5,4$, respectively. Correspondingly, they increase the probability of transitioning to state $s^{1}$ by $0.10,0.15,0.20,0.25,0.30$, respectively.
@@
    Local Treatment. The company plans to offer coupons whenever the customer’s recency is high (recency$>$2). We still consider five different candidate values. They uniformly decrease the rewards under states $s^{3},s^{4},s^{5}$ by $2,2.5,3,3.5,4$, respectively and increase the probability transiting to state $s^{1}$ by $0.10,0.11,0.12,0.13,0.14$, respectively.
We then run experiments for each candidate value under both cases treatment possible values, aiming to identify the optimal coupon value which are 4 and 2 . In each experiment, we run all three estimators over 1,000 independent trajectories and record their empirical mean and standard error. The experiment horizon is adjusted so that the IS estimator’s standard error is smaller than its empirical mean. In Figure 1, we present the simulation results.
@@
Unbiasedness across all estimators: The empirical means for all three estimators in both experiments demonstrate negligible bias. This aligns with our theoretical results of the consistency of all three estimators.
@@
Variance reduction by IS and PI: In both examples, we observe a significant reduction in variance—by at least half—through the application of IS and PI. This illustrates the effectiveness of IS and PI in overcoming the lower bound of A/B testing by utilizing the local structure. Moreover, PI consistently outperforms IS in both examples, confirming its capability in further variance reduction.
@@
While ATEs in both experiments are close, the local treatment case requires a substantially longer horizon for the standard error of IS estimator to decay comparing with the SST case. This discrepancy aligns with the intuition that the SST allows information sharing for more states than the local treatment case. This observation showcases that the efficacy of IS increases with the number of shared states.
8.2 A Real-World Sepsis Simulator
To evaluate our method in a more realistic setting, we employ a simulator of sepsis treatment for intensive care unit (ICU) patients (Komorowski et al. 2018, Oberst and Sontag 2019, Tran et al. 2023). Our focus is on experimenting with various combinations of treatments applicable to patients in critical conditions. The primary objective is to understand the long-term effects of these treatment strategies.
The Environment: The physiological status of patients can be captured by a underlying MDP. The state space encompasses discrete measurements of vital signs (heart rate, blood pressure, oxygen concentration, and glucose levels), categorized into levels (e.g., low, normal, high), along with records of previous treatments. The size of the state space size 720. The reward function assigns values of -1, 0, or +1 based on the number of abnormal vital signs exhibited by a patient. In the absence of treatment intervention, patient signals fluctuate according to specified transition probabilities, reflecting the dynamic nature of the health condition.
The Local Treatment: When patients enter a critical state—characterized by all signals being abnormal—three treatment options become available: antibiotics, vasopressors, and mechanical ventilation. This results in $n=8$ test arms, corresponding to the possible combinations of these treatments. The application of these interventions alters the transition probabilities of the various signals. In our experimental setting, the total number of treated states amounts to 128, representing approximately 18% of the overall state space.
The Simulation: Unlike the previous synthetic example, real-world constraints limit our ability to conduct extended experiments on one individual patient. Instead, we adopt a common practice in sepsis studies: assuming patients with similar contexts follow the same MDP. We pool observation data from a group of non-diabetic patients to estimate the average long-term effect of treatments. Our simulator generates observed trajectories of length 100 for each patient. These patients are initialized according to a state distribution learned a priori.
We run the simulations for $N=250$ patients over $100$ independent runs and then demonstrate the inference results of each estimators. We additionally compare two benchmarks, the Naive and Difference-in-Q (DQ) estimators, from Farias et al. (2022), which is originally designed for the general treatment policies on MDPs with two test arms. For the Naive estimators, we naively take the average on the rewards collected under the treatment policy and the control policy, ignoring the interference of state transitions. The DQ estimator strikes a balance between bias and variance under general and global treatment policies in MDPs. In Appendix 19.2, we generalize the DQ theory and provide the expression of DQ estimator for multiple test arms.
Refer to caption Refer to caption
Figure 2: Sepsis Simulator. Diamonds and caps represent the empirical mean and standard error respectively for each estimator. Left: Estimated ATE the best-performing arm among $n=8$ tested arms at patients $N=250$. Right: MSE of each estimator vs. $N$ the number of patients. IS, PI dominate over all experiment horizon.
Figure 2 summarizes the simulation results as well as the Mean Square Error (MSE) curve with the number of patients. We have the following observations.
Variance Reduction of IS and PI: We observe a significant reduction in variance through the application of IS and PI while still obtaining negligible bias comparing with the A/B testing estimator. These methods also enjoy fast decay of MSE compared to all other estimators. This confirms the effectiveness of IS and PI and demonstrate the power of leveraging treatment structure to obtain striking efficiency improvement. Moreover, aligning with our intuition, PI consistently outperforms IS, confirming its capability in further variance reduction with more information.
Unbiasedness of IS and PI: Although the Naive estimator has smaller variance than IS and comparable variance to PI, its bias is non-negligible. Furthermore, the Naive estimator struggles to exclude the null effect, potentially resulting in insufficient statistical power. In contrast, the means of the IS and PI estimators are centered around the true ATE.
DQ exhibits bias-variance trade-offs: DQ demonstrates a notable bias-variance trade-off when compared to A/B testing and Naive estimators. DQ achieves variance reduction at the cost of introducing small bias, resulting in a faster decay of MSE. However, DQ fails to leverage the inherent local structures of the treatments, leading to significantly higher MSE compared to IS and PI. In Section 9.2, we introduce a novel estimator that combines DQ and IS concepts, demonstrating improved performance over the original DQ formulation.
@@ 9.2 Bias-Variance Trade-off
With general treatment, the seminal work by Farias et al. (2022) proposes a DQ estimator that trades a small bias for a striking variance reduction in average reward setting. In a follow-up work (Farias et al. 2023), they extend the idea of DQ to other reward settings including discounted reward used in this paper. In particular, we adapt their DQ theory to obtain the following estimator